OpenStudy (anonymous):

Mr. Nichols wants to store 25 1/2 cups of stew in containers. Each container holds a maximum of 1 1/2 cups of stew. What is the minimum number of containers Mr. Nichols needs to hold all the stew?? A. 9 B. 17 C. 25 D. 51

OpenStudy (anonymous):

b. 17

OpenStudy (anonymous):

25 1/2 divided by 1 1/2 (change to improper fractions) 51/2 divided by 3/2 (flip what you are dividing by then multiply) 51/2 x 2/3 = 102/6 = 17